# PIPA: A Unified Evaluation Protocol for Diagnosing Interactive Planning Agents

### v1

@@ Appendix C Qualitative Examples of Agents’ Erroneous Behaviors
Table 6: Excerpt of failure cases in each PIPA metric.
@@ Appendix E Additional Discussions on Agent-Mixing Experiments
While mistral-large achieves higher task completion ($\mathcal{R}$) compared to llama-3.3-70B and qwen2.5-72B, it demonstrates substantially lower performance on PIPA metrics. Specifically, we observe a decline in task completion when substituting the intermediate understanding component with llama-3.3-70B (.358 $\rightarrow$ .238) or qwen2.5-72B (.358 $\rightarrow$ .336). Conversely, there is a marginal increase in the average PIPA scores (.605 $\rightarrow$ .606, .609);however, this slight improvement is unlikely to have a significant impact on overall user satisfaction.
@@
We find that qwen2.5-72B mixed with llama-3.3-70B for intermediate understanding yields the highest PIPA scores (.634) among all other rows, despite its relatively poor task completion rate (.272). Conversely, using mistral-large for intermediate understanding does not increase--and sometimes decreases--PIPA scores (.615), but it demonstrates the highest task completion scores (.367).
@@
As emphasized in Section 5.2, using qwen2.5-72B for intermediate understanding with llama-3.3-70B shows the highest improvement in task completion (.263 $\rightarrow$ .361) with only a minimal drop in the average PIPA score (.628 $\rightarrow$ .624). We observe similar, but not as impressive performance when we use mistral-large for intermediate understanding (task completion: .263 $\rightarrow$ .333; average PIPA: .628 $\rightarrow$ .587).
On top of these results, we conduct a supplementary experiment on mixing agents, referred to as the Best-of-N configuration in Table 7. In this setting, for each step (i.e., all states and responses), we collect the outputs of all three models and select the one that achieves the highest average performance across PIPA evaluation metrics (excluding task completion), as assessed by llama-3.3-70B judge. Although this configuration does not yield the highest performance in either task completion or the average PIPA score individually, it demonstrates a balanced performance between these two key indicators of user satisfaction.
@@ Appendix F Erroneous Patterns in User Simulators
Through our analysis in Section 5.3, we find that utilizing an LLM judge fails to identify numerous failure cases of the user simulator, necessitating reliance on human evaluators. However, such human-in-the-loop evaluations are both financially and logistically unsustainable. To mitigate this limitation, establishing a comprehensive and standardized protocol for a user simulator is essential (Aher et al., 2023; Yoon et al., 2024). We leave this to future work.

### v2

@@ Appendix C Qualitative Examples of Agents’ Erroneous Behaviors
Table 6: Excerpt of failure cases in each AURA metric.
@@ Appendix E Additional Discussions on Agent-Mixing Experiments
While mistral-large achieves higher task completion ($\mathcal{R}$) compared to llama-3.3-70B and qwen2.5-72B, it demonstrates substantially lower performance on AURA metrics. Specifically, we observe a decline in task completion when substituting the intermediate understanding component with llama-3.3-70B (.358 $\rightarrow$ .238) or qwen2.5-72B (.358 $\rightarrow$ .336). Conversely, there is a marginal increase in the average AURA scores (.605 $\rightarrow$ .606, .609);however, this slight improvement is unlikely to have a significant impact on overall user satisfaction.
@@
We find that qwen2.5-72B mixed with llama-3.3-70B for intermediate understanding yields the highest AURA scores (.634) among all other rows, despite its relatively poor task completion rate (.272). Conversely, using mistral-large for intermediate understanding does not increase--and sometimes decreases--AURA scores (.615), but it demonstrates the highest task completion scores (.367).
@@
As emphasized in Section 5.2, using qwen2.5-72B for intermediate understanding with llama-3.3-70B shows the highest improvement in task completion (.263 $\rightarrow$ .361) with only a minimal drop in the average AURA score (.628 $\rightarrow$ .624). We observe similar, but not as impressive performance when we use mistral-large for intermediate understanding (task completion: .263 $\rightarrow$ .333; average AURA: .628 $\rightarrow$ .587).
On top of these results, we conduct a supplementary experiment on mixing agents, referred to as the Best-of-N configuration in Table 7. In this setting, for each step (i.e., all states and responses), we collect the outputs of all three models and select the one that achieves the highest average performance across AURA evaluation metrics (excluding task completion), as assessed by llama-3.3-70B judge. Although this configuration does not yield the highest performance in either task completion or the average AURA score individually, it demonstrates a balanced performance between these two key indicators of user satisfaction.
@@ Appendix F Erroneous Patterns in User Simulators
Through our analysis in Section 5.3, we find that utilizing an LLM judge fails to identify numerous failure cases of the user simulator, necessitating reliance on human evaluators. However, such human-in-the-loop evaluations are both financially and logistically unsustainable. To mitigate this limitation, establishing a comprehensive and standardized protocol for a user simulator is essential [pmlr-v202-aher23a, yoon-etal-2024-evaluating]. We leave this to future work.